OpenStudy (anonymous):

The length of one leg of a right triangle is 48 ft, and the length of the hypotenuse is 50 ft. What is the length of the other leg?

jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 (48)^2 + b^2 = (50)^2 2304 + b^2 = 2500 b^2 = 2500 - 2304 b^2 = 196 b = sqrt(196) b = 14 So the length of the other leg is 14 ft